# Patterns of Patterns III

### v1

@@ Abstract.
Building on earlier installments, this paper re-examines the PLACARD pattern. We report on a series of workshops where PLACARD was used to scaffold collaborative reflection, speculative inquiry, and stimulate design pattern generation. These accounts are enriched by a comparison case: virtual workshops carried out with simple AI-based chatbots. We discuss limitations and lessons learned from both the human and multi-agent settings. We conclude by outlining a future development strategy at the intersection of AI agents, design patterns, and institutional governance.
Design Patterns, Pattern Languages, Action Reviews, Futures Studies, Causal Layered Analysis, Free Software, Peeragogy, Artificial Intelligence, Anticipation, Doughnut Economics
copyright: rightsretainedjournalyear: 2021doi: 10.XXXX/XXXXXXX.XXXXXXXconference: PLoP’25: Pattern Languages of Programs 2025; October 12–15, 2025; HILLSIDEbooktitle: PLoP’25, OCTOBER 12–15, HILLSIDE. Copyright 2025 is held by the author(s)price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Social and professional topicsccs: Software and its engineering Designing softwareccs: Software and its engineering Open source modelccs: Applied computing Operations researchccs: Computing methodologies Modeling and simulation
@@ 1. Introduction
In an era of increasing global interconnectivity, effective collaboration across cultural, geographical, technical, and disciplinary boundaries is essential. Traditional methods often falter when addressing the intricate layers of context inherent in distributed collaboration. To address these concerns, we have been exploring a re-theorisation of design patterns using the tools of social intelligence. First introduced in our paper “Patterns of Patterns: A Methodological Reflection on the Future of Design Pattern Methods” (PoP), the PLACARD pattern was tested in five U.S. and U.K. workshops as detailed in the sequel “Patterns of Patterns II” (PoPII) (Corneli et al., 2021; Corneli et al., 2023). To support that work, PLACARD was elaborated into a design pattern language. It has since been adapted to additional localized settings, engaging diverse participants in hybrid and in-person formats. These sessions leveraged PLACARD to bridge perspectives and envision future modes of collaboration in rapid prototyping sessions. We have also been interested in how PLACARD can be used at a slower pace, and in distributed contexts—for example, to scaffold learning in a long-running seminar series—or even more widely, across different spheres of engagement that we have described in this series of papers. In this paper, we establish another use case, by using PLACARD patterns with virtual artificial intelligence (AI) agents. We show that design pattern thinking is well-suited to use in developing epistemic AI, by which we understand artificial intelligence systems that model and manage uncertainty, provide explanations, engage in meta-cognition, and facilitate human-AI collaboration. By working between human and computer contexts, we are able to improve our understanding of the underlying patterns of knowledge acquisition, representation, and utilization. Generative AI tools contributed in other ways to this paper, by helping us make the first draft based on their analysis of the two initial papers in this series and our three summary reports. We then took the baton and revised the writing over many subsequent versions to arrive at the cohesive work you are now reading.
2. Patterns in Use
2.1. Background
In its basic design pattern format, PLACARD combines three complementary approaches for collective learning and design.
PAR (Project Action Review):
@@ 2.1. Background
– A reflective process assessing intentions, actions, and outcomes to learn from experience.
CLA (Causal Layered Analysis):
@@
– A futures studies technique analyzing issues across multiple layers—the familiar litany of well-known problems, the systems that underpin them, along with the prevailing worldviews, and deeper myths that constitute the root causes of these problems, along with the possibility of developing alternative perspectives.
DPL (Design Pattern Language):
@@
– A framework for capturing and sharing reusable solutions in a structured Context-Problem-Solution format (or similar), further structured by links between such elements.
By combining these approaches, PLACARD enables teams to cycle through reflection, deep analysis, and pattern creation. However, this is not necessarily a serial process or a fixed cycle. For instance, the generation, use, and analysis patterns may be ongoing. In our work a PAR is typically conducted after a period of work, to reflect on progress made, however it also serves as an intention-setting exercises might happen before the next phase begins. The nominal function of a CLA is to analyze a situation, but just as often they are used to describe how that situation can change.
The three papers in the series to date can be brought into schematic alignment with these methods.
(PoP $\approx$ CLA) We used CLA to think about the pattern literature as a whole and to retheorise its workings.
(PoPII $\approx$ DPL) We elaborated PLACARD into a pattern language.
(PoPIII $\approx$ PAR) We take a deep look at how these patterns have been used in practice, and take stock of potential next steps in the overall trajectory of research.
@@
Thinking about patterns in pattern-theoretic terms, we uncovered PLACARD’s modularity. Methods that achieve the same function as one of its components can be swapped in and out, without undermining the integrity of the whole. For example, PAR could be replaced with the US Army’s After Action Review (US Army, 1993)—which inspired its basic format in the first place—or by the Feedback Wheel devised by Janet Hurley (Real, 2008). CLA could be replaced by Triple-Loop Learning (Tosey et al., 2012) or Ecological Systems Theory (Bronfenbrenner, 2000). DPL could be replaced by Elinor Ostrom’s institutions (Ostrom, 1990), or by a catalogue of literary tropes, and so on. Each such selection comes with its own ideological orientation, and typical domain of application.
2.2. Implementation
We implemented PLACARD in multiple settings, tailoring our approach to each context:
@@ 2.2. Implementation
Our “PLACARD Workshop” (Danoff et al., 2024) was a hybrid focus group with one onsite and two remote facilitators (via Jitsi). Participants conducted meta-analysis on the conference itself using CLA and DPL to uncover underlying challenges, identify linguistic barriers and explore growth opportunities. One challenge discussed was how to be more inclusive of countries across the Asian continent. The conference organizers took a positive step by holding the 2025 event in Taiwan, but we cannot state that location decision was influenced by our Focus Group. In 2024, onsite attendees helped overcome barriers by clarifying for confused participants in Japanese. This was the only event in this paper conducted in multiple languages. A growth opportunity discussed that learning and networking does not need to stop when the conference itself ends. It can continue afterwards as attendees cross-pollinate pattern design collectively. The PAR was then applied to close the Focus Group and capture the results. During the evening conference recap ceremony there was one summary delivered in English by a paper author and another in Japanese by an attendee.
@@
A hybrid session titled “AI Future Envisioning with PLACARD” (Tedeschi et al., 2024) with one onsite and three remote facilitators. Participants employed the Transfeminist Tech Oracle card game as a creative prompt, then used CLA and PAR in tandem to explore speculative AI futures. Due to time constraints and abstraction load, some participants expressed confusion about the method structure, prompting later pattern refinement. A post-session PAR survey was used for reflection after the event.
@@
A common challenge for all of these workshops is rapid skill acquisition. We plan to continue this workshop series, for instance with a EuroPLoP 2025 workshop session. As mentioned above, this is a good time to take stock of what we’ve learned so far.
@@ 2.3. Reflexive Analysis Using the PLACARD Pattern Catalogue
To analyze the workshops beyond surface-level outcomes, we applied a reflexive methodology grounded in the PLACARD pattern catalogue, presented in PoPII. This catalogue is structured into three interlinked categories:
@@
For each workshop, in this paper we use the A/B/C structure to retrospectively assess:
@@
  • Where breakdowns in understanding or facilitation could be explained by missing or misapplied patterns, and
  • Where new patterns could be introduced to resolve to unanticipated tensions.
A brief summary is given in Section 3 of this paper. This reflexive approach helped us make sense of divergent workshop dynamics, and the possibility for further patterns that could help make the workshops more rewarding in general.  For example, Name the Frame was identified in hindsight as something that could have improved onboarding. Likewise, Harvest Before You Close was seen as a missing step in several sessions, and this pattern would address missed opportunities for extemporaneous pattern generation.
Rather than enforcing a rigid rubric, the catalogue functioned as a scaffold for learning from variation. It allowed us to compare facilitation dynamics, participant responses, and methodological adaptations across sites—and, crucially, to refine the catalogue itself through lived experience.
2.4. A Note on Notation
As above, we will use all-caps (like PLACARD) for patterns introduced in other sources, and small-caps (like Name the Frame) for patterns introduced in this paper, details of which are in the Appendix.
2.5. Empirical and Design-Based Development of New Patterns
Across the 2024 and 2025 events, new patterns did not emerge fully formed. Indeed, only one new pattern surfaced and named explicitly during the workshops:
  • Have Multiple Leaders Onsite – prompted by the hybrid setup at EuroPLoP, where limited physical facilitation led to friction. The pattern proposes distributed facilitation roles to support hybrid presence, reduce coordination load, and maintain session energy.
Other new proto patterns crystallized in our retrospective analysis. For example:
  • Name the Frame – (a further specification of CONTEXT SETTING from PoPII) this pattern emerged from repeated facilitator observations that participants struggled to orient themselves without a clear method overview.
  • Borrow the Situation – inferred from participant improvisation in choosing and reinterpreting prompts when predefined situations felt mismatched or abstract.
  • Start with Sense – (generalizing DÉRIVE COMIX from PoPII) builds on the observation that embodied play or metaphorical framing was often the key to unlocking participation
In sum, while few patterns emerged as polished artifacts within the workshop frame, the pattern language itself evolved through encounter. The workshops provided empirical grounding for evaluating utility, exposing limits, and highlighting latent needs. This iterative, design-based process parallels how pattern languages grow in living communities—not through top-down invention, but through grounded, reflective refinement.
@@ 2.6. Developing a comparison case: Working with patterns computationally
We noticed that it is possible to ask ChatGPT to run simple multi-agent “virtual” workshops. These can be compared with the in-person workshops. Below are two lightly edited example prompts that were used for setting up these simulations.
Prompt 1 Given to Generative AI Program I want you to create a multi-agent system with 15 different ChatGPT-based agents.  I want you to assign each of them the name of a distinct animal, and each of the agents will have some of the typical characteristics of the animal it is named for.  Then I want you to walk through the 30 “Peeragogical Exercises for a Seminar on Epistemic AI” [elided] in a virtual seminar.  The final exam is a dialogue between these animals, about the following prompts.  Please have as much discussion as you’d like among the agents before you give me their final team presentation answering the questions. What is epistemic AI? How does mathematics relate to this vision? What does ‘peeragogy’ mean, and how does it relate to this vision? What kind of multi-agent has created this answer? Why is this virtual seminar of historical importance?
Prompt 2 Given to Generative AI Program Remember how we did some simple multi-agent workshops before? Let’s simulate the architecture described the paper with some ChatGPT-based agents. This time, rather than basing them on animals, we are going to base the agents on a diverse group of Research Fellows in AI. Let’s only consider 10 such agents. Please include the “Ant” role from the workshop with animals, she will record what happens in which workshop session. Please also include the “Dog” role from the animal workshop, who will act as a primary reporter. The Fellows will be set the task of designing a series of workshops using the PLACARD method and associated patterns, and the outcome of the series should be a detailed plan of work for a research project on epistemic AI. You should do deep research to make sure that the discussions are plausible.
3. Patterns as Technical, Epistemic, and Reflective Infrastructure
3.1. Reframing Patterns as Sense-Making Scaffolds
@@
Patterns are often described in instrumental terms: reusable solutions to recurring problems in context. As we look into how they work, we see there’s more to it. One aspect of this is the use-versus-mention distinction: clearly, a physical alcove is not the same as Christopher Alexander’s “ALCOVES” pattern (Alexander et al., 1977). And at both levels, there’s a degree of explicitness. When does an indentation become an alcove? When does a pattern become part of a language?
The Start with Sense pattern characterizes various ways of inviting participants to begin inquiry with embodied, perceptual, or intuitive grounding. This creates conditions under which problems and possibilities become intelligible.
Across the A/B/C catalogue, many patterns introduced earlier and from the Appendix operate similarly:
  • Name the Frame helps surface the underlying structure of an activity or method.
  • Borrow the Situation allows participants to reinterpret or co-create the context of inquiry.
  • Meta-Reflection Loop supports collective understanding of what just happened and why.
These are cognitive footholds, ways of managing ambiguity and distributing attention. They help groups compose a shared reality, especially when dealing with complex or unfamiliar material.
From this perspective, the value of a pattern lies not just in its repeatability but in its sensitivity to situated meaning-making. A good pattern doesn’t just say “do this”; it says “look here, attend to this tension, notice this affordance.” It invites participants to see the world in terms of actionable differences.
Reframing patterns this way brings PLACARD into closer dialogue with traditions such as:
  • Phenomenology—emphasis on situated perception (Moran, 2000)
  • Participatory design—emphasis on co-articulation of problems (Spinuzzi, 2005)
  • Grounded theory—emphasis on emergent categorization (Glaser, 1992)
It also prepares the ground for computational application. For artificial agents, much as for humans, sense-making begins with activation of relevance—the ability to determine what matters in a given context. In PLACARD, patterns serve as activation scaffolds: semi-formal cues that guide what should be noticed, named, or negotiated. Each of its layers—A/B/C—corresponds to a distinct mode of learning. Crucially, design patterns—or, indeed, any substitute schema that plays the same role—achieve the function of linking these modes, enabling transitions between experience, inquiry, and abstraction.
We propose that patterns support at least three interdependent forms of learning:
  1. (1)
    Reflective Learning (Boyd and Fales, 1983): Patterns such as PROJECT ACTION REVIEW  and Meta-Reflection Loop support meaning-making after action. They help participants surface experiences, identify what shifted, and consider next steps. This mode is retrospective and grounded in collective memory and sense-making.
  2. (2)
    Design Learning (Burdick, 2012): Patterns like Borrow the Situation or Pattern the Play create scaffolds for reimagining problems, proposing alternatives, and testing ideas. Design learning is generative, situated in ambiguity. Patterns here serve as framing devices that guide creativity while constraining chaos.
  3. (3)
    Model Learning: Patterns like Name the Frame and Harvest Before You Close deploy abstraction—capturing structural insights from activity. This form of learning supports knowledge representation and the translation of local insights into portable conceptual tools (e.g., (Ariyaratne, 1977)).
Patterns enable movement between these modes. For example:
  • A reflection may generate a design insight that prompts reframing.
  • A design intervention may require formal modeling to share or extend.
  • A model may provoke a return to embodied or emotional grounding.
3.2. Summary of the evidence base from our in-person workshops
Table LABEL:workshop-summary summarizes the use of PoPII patterns in our workshops. Further details are the Appendix.
Table 1. Summary of patterns used in workshops and potential for improvement
Category/Patterns Used Well Needs Refinement
A./DÉRIVE COMIX, DO YOUR RESEARCH, CONTEXT SETTING, PAR Overrun — PAR provides shared cohesion, when used; our PAR-inspired survey has a different feel. Context-setting could also be more cohesive: Name the Frame. Explore more creative, embodied PAR.
B./GOING META, FACILITATOR and FUNCTIONAL ROLES Mixed — time + onboarding were limits. Introduce meta-language tools earlier and more clearly. Test whether embodied cues help people get to worldview/myth.
C./REINFUSE EXPERTISE Implicit — patterns used, not made. Stronger follow-through on actionable steps; Harvest Before You Close. Explicitly support pattern emergence from cards/music activities.

### v2

@@ Abstract.
Through a series of workshops, we looked at ways to structure and scaffold group dialogue, and support the emergence of novel design patterns. We contrast these sessions—which we ran with other humans—with two “virtual workshops” which we simulated with ChatGPT. Limitations in both human and virtual settings are discussed, alongside lessons learned. We conclude by proposing a development trajectory that combines AI agents, pattern-based design, and institutional governance.
Design Patterns, Pattern Languages, Artificial Intelligence
copyright: rightsretainedjournalyear: 2025doi: 10.XXXX/XXXXXXX.XXXXXXXconference: PLoP’25: Pattern Languages of Programs 2025; October 12–15, 2025; HILLSIDEbooktitle: PLoP’25, OCTOBER 12–15, HILLSIDE. Copyright 2025 is held by the author(s)price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Social and professional topicsccs: Software and its engineering Designing softwareccs: Software and its engineering Open source modelccs: Applied computing Operations researchccs: Computing methodologies Modeling and simulation
@@ 1. Introduction
PLACARD is a design pattern for building and using design patterns (Corneli et al., 2021). The pattern was explored in five U.S. and U.K. workshops (Corneli et al., 2023), and, in the process, PLACARD as a design pattern was elaborated into a design pattern language. In this paper, we report on further workshop experiences with PLACARD, and establish a further use case: using the design pattern languages with virtual artificial intelligence (AI) agents. Structurally, the paper (§2) provides all necessary background on PLACARD; (§3) outlines our methods; (§4) presents evidence from our recent workshops; (§5) proposes a set of “Design Patterns for Pattern-Competent Agents”; (§6) describes our simple multi-agent workshop simulations; (§7) discusses the study’s limitations, lessons learned, and related work; and (§8) draws conclusions.
(a) Refer to caption
(b) Refer to caption
Figure 1. Text analytics from undetectable.ai
(a) paper body (b) appendix
From a methodological standpoint, we drew on OpenAI’s ChatGPT primarily as a research catalyst, and only secondarily as a writing aid (Figure 1(a)), with the aim of exploring the extent to which contemporary chatbot technology can work effectively with design patterns (Figure 1(b)).
2. Background
2.1. Motivation
In an earlier paper, we observed:
Patterns have previously been discussed in explicitly computational terms—however, that direction of work so far remains mostly at the level of a proposal […], with only limited discipline-specific uptake[.] Could this change, to generalize the kinds of patterns that can be computed with, and make them interoperable? (Corneli et al., 2021)
What was once a speculative trajectory can now be investigated through practical experiments—using generative models to simulate, remix, and refine design pattern languages.
2.2. Terminology and notation
We use “pattern” as a shorthand for “design pattern” unless otherwise stated. We should acknowledge that our approach to patterns has more in common with the early writing of Alexander (e.g,. Alexander and Poyner (1970)) than with the more narrowly technical interpretations found in later software engineering literature. Rather than anchoring our work to any single disciplinary tradition, we explore pattern theory from the perspective of design at large, with particular attention to the meta-patterns that shape pattern authoring and use. In the process, we aim to address a gap in Alexander’s work. He largely presupposes the generative capacity of patterns, without systematically unpacking how they work or under what conditions they foster coherence, value, or transformation.
Some of our effort will be spent on further exploration of the meaning of “design pattern”; but Leitner’s (2015) definition—“Patterns are shared as complete methodic descriptions intended for practical use by experts and non-experts”—is a good starting point.
The name “PLACARD” was introduced to denote a remix of three complementary methods for organising collective learning and design (PAR, CLA, and DPL—the acronyms will be expanded and explained just below). This anagrammatic device served as a memorable handle in the context of our earlier papers, but it is not at all transparent on first encounter, nor is it quite accurate, since it implies that the pattern can only be implemented using these specific methods, which is not really the case. A more transparent and explanatory title for the pattern itself, is Patterns Embody Intelligence. The following subsection will briefly retrace the historical development of PLACARD, and introduce “Patterns Embody Intelligence” as a generalized reworking.
As above, we will use all-caps (like PLACARD) for patterns described in other sources, and small-caps (like Patterns Embody Intelligence) for patterns introduced in this paper.
2.3. PLACARD — Patterns Embody Intelligence
The PLACARD pattern was created by aligning three established methods for organising collective intelligence with the typical “context/problem/solution” design pattern schema, as in Table 1. The idea here is that the methods can be used to operationalize the creation and use of patterns.
Table 1. PLACARD components schematized
Element Method (acronym) What it is
@@ 2.3. PLACARD — Patterns Embody Intelligence
Context Project Action Review (PAR) A reflective process assessing intentions, actions, and outcomes, useful for learning from experience.
@@
Problem Causal Layered Analysis (CLA) A futures studies technique analyzing issues across multiple layers—the familiar litany of well-known problems, the systems that underpin them, along with the prevailing worldviews, and deeper myths that constitute the root causes of these problems, along with the possibility of developing alternative perspectives.
@@
Solution Design Pattern Language (DPL) A framework for capturing and sharing reusable solutions in a structured Context-Problem-Solution format (or similar), further structured by links between such elements.
This, then, is the PLACARD pattern at its simplest: a somewhat recursive pattern for creating design patterns:
PLACARD
Context: Opertating in some context, use the PAR to understand the context.
Problem: Use the CLA to understand and frame associated problems and their interconnections.
Solution: Use DPL to elaborate the solution (returning to PAR and CLA as needed).
That said, stated in this way, PLACARD is perhaps rather less a pattern than a meta-method. Our next move was to make the specific methods it mentions into placeholders, and to emphasise instead their functional role.
PLACARD
Context: By using PAR (or another sensory method), we are able to identify recurring themes within a context.
Problem: Then, by using CLA (or another cognitive method), we are able to organize these repeating themes in a structure that exposes the underlying trends, causes, and potential terminating states.
Solution: With DPL (or another motor method) we can make what we have learned actionable.
@@
For example, PAR could be replaced with the US Army’s After Action Review (US Army, 1993)—which it itself adapts—or by the Feedback Wheel devised by Janet Hurley (Real, 2008). CLA could be replaced by Triple-Loop Learning (Tosey et al., 2012) or Ecological Systems Theory (Bronfenbrenner, 2000). DPL could be replaced by Elinor Ostrom’s institutions (Ostrom, 1990), or by a catalogue of literary tropes, and so on.
A further step is to notice that by the time get to this point PAR, CLA, and DPL are no longer really needed as part of the pattern statement; they are hole-fillers. We can properly introduce and start to use new, more general, terminology, better surfacing the underlying pattern (and retitling it) this way:
Patterns Embody Intelligence
Context: We operate in a context in which intelligence is available to us. By using a suitable sensory method, we are able to characterize a particular ‘sub-context’ of interest.
Problem: Then, by using a suitable cognitive method, we are able to organize themes and tensions to identify a ‘problem’ within that sub-context.
Solution: With a suitable motor method, we can make what we have learned actionable in a way that ‘solves’ the problem.
In earlier work, we used design patterns to further decompose and operationalize each of these phases. We developed patterns for characterizing a context by identifying themes; patterns for organizing structure and exploring tensions; and patterns for transforming insights into action.
The current paper returns to this familiar terrain with a different aim: to identify and reflect on the patterns that are necessary for working with patterns at all. Rather than asking “how might pattern work flourish?”, we now ask “what makes pattern work possible?” The pattern Patterns Embody Intelligence offers a conceptual launch point for this inquiry. We will revisit and revise it later in the paper, drawing on insights from our experiments with pattern-capable agents.
3. Methods
3.1. Designing and running in-person workshops
We discussed the basic PLACARD pattern and used it (often in its more adaptable form, PLACARD) to structure workshops in multiple settings, tailoring our approach to each. Our hope was to make PLACARD useful to a broader audience, though this was largely unsuccessful, since to our knowledge the authors of this paper are currently the only users of this design pattern. Nevertheless, the workshops gave us a core empirical dataset.
@@ 3.1. Designing and running in-person workshops
Our “PLACARD Workshop” (Danoff et al., 2024) was a hybrid focus group with one onsite and two remote facilitators (via Jitsi). Participants conducted meta-analysis on the conference itself using CLA and DPL to uncover underlying challenges, identify linguistic barriers and explore growth opportunities. One challenge discussed was how to be more inclusive of countries across the Asian continent. (The conference organizers took a positive step by holding the 2025 event in Taiwan, but we cannot state that this decision was influenced by our Focus Group.) In 2024, onsite attendees helped overcome barriers by clarifying for confused participants in Japanese. This was the only event in this paper conducted in multiple languages. A growth opportunity discussed that learning and networking does not need to stop when the conference itself ends. It can continue afterwards as attendees cross-pollinate pattern design collectively. The PAR was then applied to close the Focus Group and capture the results. During the evening conference recap ceremony there was one summary delivered in English by a paper author and another in Japanese by an attendee.
@@
A hybrid session titled “AI Future Envisioning with PLACARD” (Tedeschi et al., 2024) with one onsite and three remote facilitators. Participants employed the Transfeminist Tech Oracle card game as a creative prompt, then used CLA and PAR to explore speculative AI futures. Due to time constraints and abstraction load, some participants expressed confusion about the method structure, prompting later pattern refinement. A post-session PAR survey was used for reflection after the event.
@@
EuroPLoP 2025 (Irsee, Germany):
A common challenge for all of these workshops is rapid skill acquisition. PLACARD has a lot of moving parts. In our EuroPLoP 2025 workshop session, we used the PAR to good effect, but left the other components to one side, replacing didactic instruction with more embodied play.
@@ 3.2. Reflexive Analysis Using the PLACARD Pattern Catalogue
We drew on AI support to help analyze the workshops. We imported the existing PLACARD pattern catalogue into ChatGPT. In line with the description above, the catalogue is structured in three categories:
@@
We retrospectively analysed our write-ups of the above workshops to help assess:
@@
  • Where breakdowns in understanding or facilitation could be explained by missing or misapplied patterns.
This preliminary analysis was then further developed by hand; the result is presented in Section 4 of this paper. We also considered whether new patterns might be introduced that would resolve tensions that had not been anticipated in PoPII; these patterns appear in the Appendix.
@@ 3.3. Developing a comparison case: Working with patterns computationally
We noticed that it is possible to ask ChatGPT to run simple multi-agent “virtual” workshops. We ran several such simulated workshops, of which we will discuss two. Below is a generic example prompt for creating these light-weight simulations.
Prompt Template: Multi-Agent Simulation for Collaborative Inquiry Simulate a structured multi-agent workshop involving $N$ distinct agents, each based on a unique [metaphorical persona or professional identity], and endowed with characteristic traits or expertise aligned with that identity. The agents should collaborate on a shared task: designing a set of [outputs, e.g., workshops, protocols, strategies] using the [specified framework or method, e.g., PLACARD, Design Patterns, Theory of Change].
Assign the following roles:
A Recorder agent who captures the evolving structure and rationale of the discussion.
A Reporter agent who periodically summarises the current state of collective understanding.
Optionally, include other roles, such as a Socratic Facilitator agent who prompts clarification, challenges assumptions, and encourages epistemic humility.
The agents should engage in $M$ discrete sessions, with each session addressing a defined phase or subtask. The dialogue should be plausible, reflective of diverse perspectives, and demonstrate progression over time.
After all sessions are complete, generate a synthetic report or artefact that includes:
The main findings or decisions of the group
A reflection on the group process and emergent dynamics
A meta-level commentary on the kind of multi-agent system that was instantiated
Please ensure internal consistency and draw on relevant external knowledge or frameworks where applicable.
4. Patterns as Reflective Infrastructure
@@
Patterns are often described in instrumental terms: reusable solutions to recurring problems in context. However, there’s clearly more to it than that. One aspect of this is the use-versus-mention distinction, for example, a physical alcove is not the same as the “ALCOVES” pattern (Alexander et al., 1977). There’s also a degree of explicitness. When does an indentation become an alcove? When does a proto-pattern become a pattern? One typical answer to the latter question is that the pattern should be applicable in suitably varied contexts. A further implication is that a pattern might be “observed” empirically even when it has not been formally “applied” in a design sense. After all, alcoves existed long before 1977!
Although we did not explicitly structure our recent series of workshops using PLACARD pattern language (rather, we used the earlier PLACARD pattern), nevertheless the workshops continue the spirit of that work, and the patterns themselves could often be rationally reconstructed from the evidence base that the current round of workshops provides. Details of the previous pattern statements are to be found in (Corneli et al., 2023). The qualitative analysis here compiles evidence of their implicit use, or the use of a similar strategy, looking across the in-person workshops mentioned above. This is not intended to be a ‘feature grid’ showing which workshops used which patterns.
A. Identifying Themes
DÉRIVE COMIX
Application: In several of the workshops, participants used Transfeminist Tech Oracle cards (with ‘suits’: object, value, body/territory, situation) as interpretive prompts to spark imagination—unfamiliar juxtapositions mirroring the urban dérive or imagined stroll in a more conceptual sense. Observation: The card game facilitated divergent thinking akin to exploring unfamiliar terrain. In the 2025 EuroPLoP workshop, we did a similar exploration, using less complex images from the card game Dixit.
CONTEXT SETTING
Application: Early slides and explanations introduced key concepts (e.g., design patterns, CLA). Observation: Participant feedback highlighted a need for clearer up-front orientation. Improved alignment with this pattern could increase accessibility. That said, in the 2025 EuroPLoP workshop, observed that some “interface friction” can sometimes be a benefit for workshop outcomes, even if it is frustrating for participants.
SHARE BACK
Application: When there were a sufficient number of attendees, they did break into small groups and share back.
DO YOUR RESEARCH
Application: Past workshop experiences were integrated into the design; earlier in-house uses of CLA and PAR informed subsequent sessions. Example: Reference to Anticipation 2019 and prior events illustrates “pre-research” informing present design.
B. Organizing Structure
GOING META
Application: The overall project itself exemplifies recursive methodology—using PLACARD to better understand and refine PLACARD. Example: In two of the workshops, participants were prompted to reflect on and imagine both the methodology and the subject (AI futures).
PATTERN LANGUAGE COMPONENTS
Weak application: The workshops in this series typically focused on the PLACARD components (CLA + DPL + PAR). We typically did not go into further detail of DPLs. While it is true that CLA, DPL, and PAR do suggest a way to think about pattern language components, because they are complex, they are not always user friendly. Recommendation: In some of the workshops reported on in our previous paper, we used light-weight keywords (“however”, “because”, “specifically”) to scaffold understanding of design pattern internals incrementally. In the current series of workshops, that might have been less useful in any case, because attendees were already familiar with pattern languages. We had supposed that, within this population, the Transfeminist Tech Oracle cards would prompt pattern-like thinking, but we did not assess whether it did or not.
FACILITATOR ROLES
Application: Hybrid facilitation with distinct roles (onsite leader, remote supporters) distributed facilitation and modelled different levels of participation. (Cf. FUNCTIONAL ROLES below.)
C. Making it Actionable
THE FUTURE BEGINS NOW
Weak application: Using the card game, participants generated imagined futures. Recommendation: While creative ideas about the future emerged, follow-through actions or small-scale pilots were typically not described. In the 2025 EuroPLoP workshop, and other contexts where a PAR was carried out, “next steps” were explicitly surfaced. These were often next steps for the workshop leaders, not workshop participants, but in the 2025 EuroPLoP session, we explicitly directed participants to think about potential next steps for themselves.
FUNCTIONAL ROLES
Weak application: We did not explicitly involve workshop participants in any formal roles. (In the 2025 EuroPLoP workshop, participants did have the opportunity to take on an untitled light-weight roles, such as introducing a simple musical theme in a circle music exercise.) Recommendation: Our earlier work developed several explicit roles for participants to explore: these or other roles could be explained and offered, even if they are not taken up—however, we must acknowledge that there is a time-cost and cognitive load when doing so.
PROJECT ACTION REVIEW (PAR)
Application: Several sessions used the PAR pattern at the end to reflect, share learnings, and gather insights. In some cases where the PAR was not used, we shared a survey with questions informed by the PAR. Recommendation: An in-person PAR and a survey are very different experiences for participants.
REINFUSE EXPERTISE
Application: Participants brought diverse professional and cultural perspectives. Their expertise was implicitly integrated into the discussion in shared reflection on topics such as AI futures and governance. Recommendation: A more explicit arc—temporarily setting aside expertise to foster common ground—is more in the spirit of this pattern. In the 2025 EuroPLoP session, participant expertise was explicitly invited in the concluding PAR.